# Av. Diagonal Las Torres, 2640 Peñalolen, Chile

### v1

@@ 3.2 Energy functionals coming from LPP CG
Extending the prescription of the previous section to six dimensions, we evaluate the invariants $I_{1}$, $I_{2}$ and $I_{3}$ in the orbifold. However, the situation is more delicate than in four dimensions. In general, cubic curvature invariants are sensitive to the so-called splitting problem Miao:2014nxa ; Camps:2014voa ; Miao:2015iba . This means that there are different naive ways to regularize the action near the conical singularity and, thus, different holographic EE functionals are obtained depending on the regularization scheme. [^10] There are higher-curvature theories for which this issue does not arise, such as quadratic gravity Fursaev:2013fta ; Anastasiou:2021swo , $f(R)$ gravity or Lovelock gravity Camps:2013zua ; Hung:2011xb ; deBoer:2011wk ; Anastasiou:2021jcv . This also occurs when the coupling constants of any higher-curvature term are treated perturbatively Bueno:2020uxs . However, it was shown in Ref. Miao:2015iba that bulk conformal symmetry should be induced to the resulting codimension-two functional. This allows to parametrize a family of splittings where the universal part of the holographic EE is independent of the specific choice. Exploiting this remarkable property of combination (52), one ends up with the following expression Miao:2015iba
@@
$+6\varepsilon^{\alpha\beta}\varepsilon^{\gamma\delta}K_{\alpha}{}^{\langle\zeta\iota\rangle}K_{\beta}{}^{\langle\eta}{}_{\iota\rangle}\left(2W_{\gamma\eta\delta\zeta}+K_{\delta\langle\eta\lambda}K_{\delta}{}^{\langle\lambda}{}_{\zeta}\right)\,,$
@@
For $F_{1}$ and $F_{2}$, we chose to maintain the covariant formulation in order to simplify the resulting expressions, considering $\varepsilon_{\alpha\beta}=n_{\alpha}^{A}n_{\beta}^{B}\epsilon_{AB}$ as the binormal, where $\epsilon_{AB}$ is the Levi-Civita tensor, and identifying $g^{\perp}_{\alpha\beta}$ as the two-dimensional metric of the normal bundle. For later convenience, we perform the normal decomposition of the $F_{3}$, by labeling the bulk indices as $\alpha=\left(A,a\right)$, where $a$ denotes tangential indices and $A$, the normal directions to $\Sigma$. Here, $C_{\alpha\beta\gamma}$ and $B_{\alpha\beta}$ represent the Cotton and Bach tensor of $\mathcal{M}$, respectively, where $\nabla_{a}^{\Sigma}$ is the covariant derivative with respect to the codimension-two intrinsic metric $\gamma_{ab}$. Further details on the normal decomposition of $F_{3}$ is given in Appendix C.
@@
This exact form will allow us to make contact with higher-dimensional analogues of both the reduced Hawking mass and Willmore energy. However, the computation of the former strongly depends on the determination of the boundary term in Eq. (65), which is a quite challenging task due to he presence of covariant derivatives that makes the computation of their conical part rather complicated. Interestingly, the form of the boundary term will be greatly simplified when restricting ourselves to Einstein spacetimes. The four-dimensional analysis in Sec. 2 indicates that both functionals arise when considering surfaces embedded in this exact class of spacetimes.
@@ 3.3 Generalized reduced Hawking mass and renormalized area
Obtaining a candidate for the reduced Hawking mass of a four-dimensional hypersurface embedded in a six-dimensional bulk spacetime is a highly non-trivial task, since this has to meet the same criteria as its two-dimensional analogue, but for a functional constructed out of four-derivative objects. As it was reveiled in Ref. Anastasiou:2022ljq , and discussed in the previous section, the two-dimensional reduced Hawking mass arises from the Graham-Witten anomaly functional $\mathbf{L}(\Sigma)$ when evaluated for Einstein spacetimes (28). Similarly, we present a candidate for the four-dimensional reduced Hawking mass, as the functional coming from the evaluation of the Graham-Reichert anomaly $\mathbf{F}(\Sigma)$ in Einstein spacetimes. In this case, the missing boundary contribution in Eq. (65) — which we denoted b. t. — can be determined, as it comes entirely from the $I_{3}$ conformal invariant. Indeed, the surviving boundary contribution from the CG action (57) when evaluated in the Einstein sector is
@@
$F(A)=\frac{\mathbf{A}^{\text{ren}}({\Sigma_{A}})}{4G_{\text{N}}}=-\frac{1}{4G_{\text{N}}}\mathbf{F}({\Sigma_{A}})\big{|}_{\text{E}}\,.$ (72)
@@ 3.4 Generalized Willmore energy
$\mathcal{J}_{{\Sigma_{\text{cl}}}}=\frac{1}{48}\left[(\partial\tilde{K})^{2}-\tilde{K}\tilde{K}^{ab}\tilde{K}\tilde{K}_{ab}+\frac{7}{16}\tilde{K}^{4}\right]\,.$ (75)
This expression coincides with the generalization of the Willmore energy for closed four-dimensional surfaces given previously in the mathematical literature [^13] Note that in the derivation presented in Ref. guven2005conformally , a factor of $-2$ is dropped in the calculations, obtaining in the end incorrect coefficients for $\partial_{a}\tilde{K}^{A}\partial^{a}\tilde{K}_{A}$ and $\tilde{K}^{A}\tilde{K}_{A}{}^{ab}\tilde{K}_{B}\tilde{K}^{B}{}_{ab}$. guven2005conformally ; Gover:2016buc ; zhang2017grahamwittens ; Graham:2017bew .
@@
This equation is one of the main results of our work, so we devote next subsection to test it for different entangling regions, namely, the round ball, small deformations of it and thin strips.
@@ 3.5.1 Sphere
The setup can be seen in Figure 2. Using this, it is an straightforward exercise to check that the bare area $\mathbf{A}\left({\Sigma_{A}}\right)=\int_{{\Sigma_{A}}}\text{d}^{4}y\sqrt{\gamma}$ of the codimension-two RT surface, equipped with metric $\gamma_{ab}$, is given by Ryu:2006ef ; Nishioka:2009un
@@
Figure 1: Ball-shaped entangling region $A$ with radius $\rho=R$ and its cobordant $\partial A=\partial\Sigma$ codimension-two surface $\Sigma$.
@@ 3.5.3 Infinite strip
Figure 3: Infinite strip entangling region $A$ of width $l$ in the $u$ direction. We introduced $L_{i}$ as IR regulators for the transversal directions $x_{i}$. The surface $\Sigma$ is cobordant $\partial A=\partial\Sigma$, where $\partial\Sigma$ composed by two parallel boundaries located at $u=\pm l/2$, this is $\partial\Sigma=\partial\Sigma^{l/2}\cup\partial\Sigma^{-l/2}$.
Figure 4: Double-copied surfaces $\Sigma$ and $\Sigma^{\prime}$ glued along the umbilical line defined by $\partial\Sigma_{1}=\partial\Sigma_{1}^{\prime}$ located at $u=\pm l/2$. The introduction of the IR regulators $L_{i}$ defines a pair of symmetric boundaries $\partial\Sigma_{2}$ and $\partial\Sigma_{2}^{\prime}$ pointing upwards and downwards along the transversal directions $x_{i}$.
@@ 4 $F(A)$ has no global bounds for $d=5$ CFTs
Figure 5: Linear-logarithmic plot of $\mathbf{W}_{5}\left(\Sigma_{a}^{(i)}\right)/\mathbf{W}_{5}\left(\mathbb{S}^{4}\right)$ with $i=1,2$, representing the $(1,1,1,1,a)$ (red line) and $(1,1,1,a,a)$ (blue line) ellipsoids respectively for different values of $a$. The values represented for the former are analytical while the latter are numerical. The red region for small values of $a$ indicates that the $(1,1,1,a,a)$ ellipsoid oscillates wildly between $\infty$- ∞ and +$\infty$+ ∞ in this regime. We also include the point $a=1$ corresponding to a round sphere $\mathbb{S}^{4}$ and represented by an eight-pointed star. As we can see, while $\mathbb{S}^{4}$ — corresponding to a $A=\mathbb{B}^{4}$ holographic entangling region — is a local minimum of the generalized Willmore functional, this is neither bounded from below nor from above for general regions.
@@ Appendix A Notation and conventions
It is important to note that in the $d+2$ decomposition that we perform here, the normal bundle indices play the role of labels for objects residing in the hypersurface. This becomes evident when the Gauss-Codazzi relations for the Christoffel symbol are considered. In particular, these can be cast as
@@ Appendix B Conformal covariantization of Einstein-AdS gravity
where we denoted $\hat{J}^{\alpha}\equiv 8W^{\alpha\gamma\delta\beta}C_{\gamma\lambda\beta}-W^{\gamma\delta}_{\beta\varepsilon}\nabla^{\alpha}W^{\beta\varepsilon}_{\gamma\delta}$. As a consequence, the scalar density $\mathcal{I}_{4}$ is Weyl invariant, i.e. $\updelta_{\varphi}\mathcal{I}_{4}=0$, where
@@ D.2 Small deformation of the sphere
Finally, we need the partial trace of the Weyl tensor at the conformal boundary and the quadratic contraction of traceless extrinsic curvature of $\partial\Sigma$ embedded in $\partial\mathcal{M}$. The first one vanishes identically for our metric under consideration $w_{ij}^{ij}=0$ whereas the second reads
@@ D.3 Infinite strip
The conformal boundary of this geometry is located at $z=\delta$, and the induced metric at this locus is just flat space with a conformal factor, i.e.
@@
where we used that for the infinite strip we have two parallel boundaries, one located at $u=l/2$, say ${\Sigma_{A}}^{l/2}$, and another symmetric one located at $u=-l/2$, i.e. ${\Sigma_{A}}^{-l/2}$ —see Figure 4. Thus, in the boundary term, there is a factor of two as they contribute symmetrically $\partial{\Sigma_{A}}=\partial{\Sigma_{A}}^{l/2}\cup\partial{\Sigma_{A}}^{-l/2}$. With these considerations, we observe that the additional term appearing in Eq. (167) cancels the divergence coming from the bare area of the RT surface (103).

### v2

@@ 3.2 Energy functionals coming from LPP CG
Extending the prescription of the previous section to six dimensions, we evaluate the invariants $I_{1}$, $I_{2}$ and $I_{3}$ in the orbifold. However, the situation is more delicate than in four dimensions. In general, cubic curvature invariants are sensitive to the so-called splitting problem Miao:2014nxa ; Camps:2014voa ; Miao:2015iba . This means that there are different naive ways to regularize the action near the conical singularity and, thus, different holographic EE functionals are obtained depending on the regularization scheme. [^10] There are higher-curvature theories for which this issue does not arise, such as quadratic gravity Fursaev:2013fta ; Anastasiou:2021swo , $f(R)$ gravity or Lovelock gravity Camps:2013zua ; Hung:2011xb ; deBoer:2011wk ; Anastasiou:2021jcv . This also occurs when the coupling constants of any higher-curvature term are treated perturbatively Bueno:2020uxs . However, it was shown in Ref. Miao:2015iba that bulk conformal symmetry should be induced to the resulting codimension-two functional. This allows to parametrize a family of splittings where the universal part of the holographic EE is independent of the specific choice. Exploiting this remarkable property of the combination (52), one ends up with the following expression Miao:2015iba
@@
$+6\varepsilon^{\alpha\beta}\varepsilon^{\gamma\delta}K_{\alpha}{}^{\langle\zeta\iota\rangle}K_{\beta}{}^{\langle\eta}{}_{\iota\rangle}\left(2W_{\gamma\eta\delta\zeta}+K_{\delta\langle\eta\lambda\rangle}K_{\delta}{}^{\langle\lambda}{}_{\zeta\rangle}\right)\,,$
@@
$\left.\left.+\nabla_{a}^{\Sigma}\left(K^{A}S_{Ab}\right)-K^{A}K^{B}{}_{ab}S_{AB}\right)\right]\,,$ (62)
and $K^{\gamma}{}_{\alpha\beta}$ is defined in Eq. (115). For $F_{1}$ and $F_{2}$, we chose to maintain the covariant formulation in order to simplify the resulting expressions, considering $\varepsilon_{\alpha\beta}=n_{\alpha}^{A}n_{\beta}^{B}\epsilon_{AB}$ as the binormal, where $\epsilon_{AB}$ is the Levi-Civita tensor, and identifying $g^{\perp}_{\alpha\beta}$ as the two-dimensional metric of the normal bundle. For later convenience, we perform the normal decomposition of the $F_{3}$, by labeling the bulk indices as $\alpha=\left(A,a\right)$, where $a$ denotes tangential indices and $A$, the normal directions to $\Sigma$. Here, $C_{\alpha\beta\gamma}$ and $B_{\alpha\beta}$ represent the Cotton and Bach tensor of $\mathcal{M}$, respectively, where $\nabla_{a}^{\Sigma}$ is the covariant derivative with respect to the codimension-two intrinsic metric $\gamma_{ab}$. Further details on the normal decomposition of $F_{3}$ are given in Appendix C.
@@
This exact form will allow us to make contact with higher-dimensional analogues of both the reduced Hawking mass and Willmore energy. However, the computation of the former strongly depends on the determination of the boundary term in Eq. (65), which is a quite challenging task due to he presence of covariant derivatives, what makes the computation of their conical part rather complicated. Interestingly, the form of the boundary term will be greatly simplified when restricting ourselves to Einstein spacetimes. The four-dimensional analysis in Sec. 2 indicates that both functionals arise when considering surfaces embedded in this exact class of spacetimes.
@@ 3.3 Generalized reduced Hawking mass and renormalized area
Obtaining a candidate for the reduced Hawking mass of a four-dimensional hypersurface embedded in a six-dimensional bulk spacetime is a highly non-trivial task, since this has to meet the same criteria as its two-dimensional analogue, but for a functional constructed out of four-derivative objects. As it was revealed in Ref. Anastasiou:2022ljq , and discussed in the previous section, the two-dimensional reduced Hawking mass arises from the Graham-Witten anomaly functional $\mathbf{L}(\Sigma)$ when evaluated for Einstein spacetimes (28). Similarly, we present a candidate for the four-dimensional reduced Hawking mass, as the functional coming from the evaluation of the Graham-Reichert anomaly $\mathbf{F}(\Sigma)$ in Einstein spacetimes. In this case, the missing boundary contribution in Eq. (65) — which we denoted b. t. — can be determined, as it comes entirely from the $I_{3}$ conformal invariant. Indeed, the surviving boundary contribution from the CG action (57) when evaluated in the Einstein sector is
@@
$F(A)=\frac{\mathbf{A}^{\text{ren}}({\Sigma_{A}})}{4G_{\text{N}}}=\frac{1}{4G_{\text{N}}}\mathbf{F}({\Sigma_{A}})\big{|}_{\text{E}}\,.$ (72)
@@ 3.4 Generalized Willmore energy
$\mathcal{J}_{{\Sigma_{\text{cl}}}}=\frac{1}{48}\left[(\partial\tilde{K})^{2}-\tilde{K}\tilde{K}^{ab}\tilde{K}\tilde{K}_{ab}+\frac{7}{16}\tilde{K}^{4}\right]\,,$ (75)
where the tildes indicate that the quantities are evaluated in $\mathbb{R}^{5}$ as ambient space. This expression coincides with the generalization of the Willmore energy for closed four-dimensional surfaces given previously in the mathematical literature [^13] Note that in the derivation presented in Ref. guven2005conformally , a factor of $-2$ is dropped in the calculations, obtaining in the end incorrect coefficients for $\partial_{a}\tilde{K}^{A}\partial^{a}\tilde{K}_{A}$ and $\tilde{K}^{A}\tilde{K}_{A}{}^{ab}\tilde{K}_{B}\tilde{K}^{B}{}_{ab}$. guven2005conformally ; Gover:2016buc ; zhang2017grahamwittens ; Graham:2017bew ; Blitz:2021qbp .
@@
This equation is one of the main results of our work, so we devote the next subsection to test it for different entangling regions, namely, the round ball, small deformations of it and thin strips.
@@ 3.5.1 Sphere
The setup can be seen in Figure 2. Using this, it is a straightforward exercise to check that the bare area $\mathbf{A}\left({\Sigma_{A}}\right)=\int_{{\Sigma_{A}}}\text{d}^{4}y\sqrt{\gamma}$ of the codimension-two RT surface, equipped with metric $\gamma_{ab}$, is given by Ryu:2006ef ; Nishioka:2009un
@@
Figure 1: Ball-shaped entangling region $A$ with radius $\rho=R$ and its cobordant ($\partial A=\partial\Sigma$) codimension-two surface $\Sigma$.
@@ 3.5.3 Infinite strip
Figure 3: Infinite strip entangling region $A$ of width $l$ in the $u$ direction. We introduced $L_{i}$ as IR regulators for the transversal directions $x_{i}$. The surface $\Sigma$ is cobordant ($\partial A=\partial\Sigma$), where $\partial\Sigma$ is composed by two parallel boundaries located at $u=\pm l/2$, i.e., $\partial\Sigma=\partial\Sigma^{l/2}\cup\partial\Sigma^{-l/2}$.
Figure 4: Double-copied surfaces $\Sigma$ and $\Sigma^{\prime}$ glued along the umbilical line defined by $\partial\Sigma_{1}=\partial\Sigma_{1}^{\prime}$ located at $u=\pm l/2$. The introduction of the IR regulators $L_{i}$ defines a pair of symmetric boundaries $\partial\Sigma_{2}$ and $\partial\Sigma_{2}^{\prime}$ pointing upwards and downwards along the transverse directions $x_{i}$.
@@ 4 $F(A)$ has no global bounds for $d=5$ CFTs
Figure 5: Linear-logarithmic plot of $\mathbf{W}_{5}\left(\Sigma_{a}^{(i)}\right)/\mathbf{W}_{5}\left(\mathbb{S}^{4}\right)$ with $i=1,2$, representing the $(1,1,1,1,a)$ (blue line) and $(1,1,1,a,a)$ (red line) ellipsoids respectively for different values of $a$. The values represented for the former are analytical while the latter are numerical. The red region for small values of $a$ indicates that the $(1,1,1,a,a)$ ellipsoid oscillates wildly between $\infty$- ∞ and +$\infty$+ ∞ in this regime. We also include the point $a=1$ corresponding to a round sphere $\mathbb{S}^{4}$ and represented by an eight-pointed star. As we can see, while $\mathbb{S}^{4}$ — corresponding to an $A=\mathbb{B}^{4}$ holographic entangling region — is a local minimum of the generalized Willmore functional, this is neither bounded from below nor from above for general regions.
@@ Appendix A Notation and conventions
In addition, we can write the contraction of $K^{A}{}_{ab}$ with bulk indices as
$K^{\gamma}{}_{\alpha\beta}=\gamma_{\alpha}^{a}\gamma_{\beta}^{b}N_{A}{}^{\gamma}K^{A}{}_{ab}\,.$ (115)
It is important to note that in the codimension-two decomposition that we perform here, the normal bundle indices play the role of labels for objects residing in the hypersurface. This becomes evident when the Gauss-Codazzi relations for the Christoffel symbol are considered. In particular, these can be cast as
@@ Appendix B Conformal covariantization of Einstein-AdS gravity
where we denoted $\hat{J}^{\alpha}\equiv 8W^{\alpha\gamma\delta\beta}C_{\gamma\lambda\beta}-W^{\gamma\delta}_{\beta\varepsilon}\nabla^{\alpha}W^{\beta\varepsilon}_{\gamma\delta}$. As a consequence, the scalar density $\mathcal{I}_{4}$ is Weyl invariant, i.e., $\updelta_{\varphi}\mathcal{I}_{4}=0$, where
@@ D.2 Small deformation of the sphere
Finally, we need the partial trace of the Weyl tensor at the conformal boundary and the quadratic contraction of traceless extrinsic curvature of $\partial\Sigma$ embedded in $\partial\mathcal{M}$. The first one vanishes identically for our metric under consideration ($w_{ij}^{ij}=0$) whereas the second reads
@@ D.3 Infinite strip
The conformal boundary of this geometry is located at $z=\delta$, and the induced metric at this locus is just flat space with a conformal factor, i.e.,
@@
where we used that for the infinite strip we have two parallel boundaries, one located at $u=l/2$, say ${\Sigma_{A}}^{l/2}$, and another symmetric one located at $u=-l/2$, i.e., ${\Sigma_{A}}^{-l/2}$ —see Figure 4. Thus, in the boundary term, there is a factor of two as they contribute symmetrically, i.e., $\partial{\Sigma_{A}}=\partial{\Sigma_{A}}^{l/2}\cup\partial{\Sigma_{A}}^{-l/2}$. With these considerations, we observe that the additional term appearing in Eq. (168) cancels the divergence coming from the bare area of the RT surface (103).